# FLAP: Fully-controllable Audio-driven Portrait Video Generation through 3D head conditioned diffusion model

### v2

@@ Abstract
Diffusion-based video generation techniques have significantly improved zero-shot talking-head avatar generation, enhancing the naturalness of both head motion and facial expressions. However, existing methods suffer from poor controllability, making them less applicable to real-world scenarios such as filmmaking and live streaming for e-commerce. To address this limitation, we propose FLAP, a novel approach that integrates explicit 3D intermediate parameters (head poses and facial expressions) into the diffusion model for end-to-end generation of realistic portrait videos. The proposed architecture allows the model to generate vivid portrait videos from audio while simultaneously incorporating additional control signals, such as head rotation angles and eye-blinking frequency. Furthermore, the decoupling of head pose and facial expression allows for independent control of each, offering precise manipulation of both the avatar’s pose and facial expressions. We also demonstrate its flexibility in integrating with existing 3D head generation methods, bridging the gap between 3D model-based approaches and end-to-end diffusion techniques. Extensive experiments show that our method outperforms recent audio-driven portrait video models in both naturalness and controllability. The source code and models are available at (URL omitted for double-blind reviewing).
@@ 4.2 Controlling Diffusion Model with 3D Head Coefficient as Condition
MCNet video - - 0.4001 7.4600 10.3600 0.2992
EchoMimic ★ ★ video - - 0.5890 6.7896 9.7304 0.3919
EDTalk ★ ★ ★ video - - 0.4159 8.4260 9.2331 0.3018
X-Portrait video - - 0.6544 5.8181 10.2428 0.4795
LivePortrait video - - 0.5474 7.6228 10.4424 0.3987
FLAP (ours) ★ ★ ★ ★ video - - 0.6107 7.7837 9.8302 0.4644
SadTalker audio 163.31 551.94 0.4798 6.9274 8.1210 0.3137
EDTalk ★ ★ ★ audio 166.11 592.20 0.4434 7.2472 8.1161 0.3052
EchoMimic ★ ★ audio 148.91 500.56 0.6116 7.0341 8.6307 0.4329
AniPortrait audio 137.27 581.01 0.6518 7.6117 8.3738 0.4920
Hallo audio 139.54 457.43 0.610 7.5069 8.1786 0.4417
Hallo2 audio 141.54 483.95 0.5993 7.7184 8.0047 0.4241
FLAP (ours) ★ ★ ★ ★ audio 136.54 399.78 0.6254 7.8617 8.2916 0.4592
@@ 5.2.2 Qualitative Evaluation
For qualitative comparisons, as shown in LABEL:fig:vis-cmp, our method generates the best results in common scenarios and scenarios with wide-range of head rotations, complex headpieces and makeup. Compared to our methods, Echomimic uses projected landmark images as control signal for generation and yields animations results with inaccurate pose and artifacts in headpieces. We perceive this inaccuracy as the loss of 3D pose information when projecting 3D key points to 2D images. As for X-portrait, it generate texture with high quality which results in better IQA scores, but it suffers from severe distortion around face and neck. Among all these methods, our method generate portrait videos with superior perceptual quality, motion precision and less artifact in both common scenarios and more complex scenarios.
@@ 5.4 Comparison with Landmark-based methods
Prior studies like Aniportrait and Echomimic use facial landmarks as guidance for controllable generation. Here we provide a more detailed comparison between our method and Aniportrait. For Echomimic, visual comparison can be found in LABEL:fig:vis-cmp. As shown in Figure 5, Lmk-based methods refers to Aniportrait. It is not capable of generating head movement like spinning and nodding since the projected 2D landmarks have lost most of the 3D pose information, thus offering vague pose information in denoising process. In Figure 5, the head generated by Aniportait in the first 3 columns from left are all squished due to this projection error. In the remaining columns, aniportrait failed to generate accurate facial dynamics when there are significant differences between the driving identity and reference identity. As we disscussed in former sections, these are the natural shortcomings of landmarks serving as control signals. However, in audio-driven mode, Aniportrait hardly generate head movement or dramatic facial expressions from audio, making it easier to maintain visual quality than other methods. On the other hand, our model use unified FLAME coefficients as conditions, avoiding the need for alignment and thus generating accurate facial dynamics and head pose as shown in Figure 5. Our approach demonstrates that FLAME coefficients can capture more subtle expressions, and is a more suitable as intermediate representation for controllable generation of talking head avatars.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A Appendix
Figure 1: Additional Results on Multi Pose Generation

### v3

@@ Abstract
Diffusion-based video generation techniques have significantly improved zero-shot talking-head avatar generation, enhancing the naturalness of both head motion and facial expressions. However, existing methods suffer from poor controllability, making them less applicable to real-world scenarios such as filmmaking and live streaming for e-commerce. To address this limitation, we propose FLAP, a novel approach that integrates explicit 3D intermediate parameters (head poses and facial expressions) into the diffusion model for end-to-end generation of realistic portrait videos. The proposed architecture allows the model to generate vivid portrait videos from audio while simultaneously incorporating additional control signals, such as head rotation angles and eye-blinking frequency. Furthermore, the decoupling of head pose and facial expression allows for independent control of each, offering precise manipulation of both the avatar’s pose and facial expressions. We also demonstrate its flexibility in integrating with existing 3D head generation methods, bridging the gap between 3D model-based approaches and end-to-end diffusion techniques. Extensive experiments show that our method outperforms recent audio-driven portrait video models in both naturalness and controllability.
@@ 4.2 Controlling Diffusion Model with 3D Head Coefficient as Condition
MCNet video - - 0.4001 7.4600 0.2992
EchoMimic ★ ★ video - - 0.5890 6.7896 0.3919
EDTalk ★ ★ ★ video - - 0.4159 8.4260 0.3018
X-Portrait video - - 0.6544 5.8181 0.4795
LivePortrait video - - 0.5474 7.6228 0.3987
FLAP (ours) ★ ★ ★ ★ video - - 0.6107 7.7837 0.4644
SadTalker audio 163.31 551.94 0.4798 6.9274 0.3137
EDTalk ★ ★ ★ audio 166.11 592.20 0.4434 7.2472 0.3052
EchoMimic ★ ★ audio 148.91 500.56 0.6116 7.0341 0.4329
AniPortrait audio 137.27 581.01 0.6518 7.6117 0.4920
Hallo audio 139.54 457.43 0.610 7.5069 0.4417
Hallo2 audio 141.54 483.95 0.5993 7.7184 0.4241
FLAP (ours) ★ ★ ★ ★ audio 136.54 399.78 0.6254 7.8617 0.4592
@@ 4.4 Audio to FLAME Module
Figure 4: Visual Results on 3 dof head pose control and rich expressions. Expressions of reference images are all neutral
@@ 5.1 Implement Details
5.2 Visual Results on Controllable Generation
Figure 4 presents some results generated by our method. Given a sequence of head angles as control signal, our model can generate film-like scene with an avatar performing user designed actions. The first row demonstrate two avatars turning his/her head around and making a head-down gesture. The second row shows an avatar first looking around and then performing another sequence of user defined actions. In the third row, we demonstrate the model’s ability of generating life-like expressions for stylized characters and novel anime characters, despite our model being trained exclusively on real portrait videos. With these abilities, our model breaks down the barriers in filmmaking, enabling real contributions to the creative process.
@@ 5.3.2 Qualitative Evaluation
For qualitative comparisons, as shown in Figure 3, our method generates the best results in common scenarios and scenarios with wide-range of head rotations, complex headpieces and makeup. Compared to our methods, Echomimic uses projected landmark images as control signal for generation and yields animations results with inaccurate pose and artifacts in headpieces. We perceive this inaccuracy as the loss of 3D pose information when projecting 3D key points to 2D images. As for X-portrait, it generate texture with high quality which results in better IQA scores, but it suffers from severe distortion around face and neck. Among all these methods, our method generate portrait videos with superior perceptual quality, motion precision and less artifact in both common scenarios and more complex scenarios.
@@ 5.5 Comparison with Landmark-based methods
Prior studies like Aniportrait and Echomimic use facial landmarks as guidance for controllable generation. Here we provide a more detailed comparison between our method and Aniportrait. For Echomimic, visual comparison can be found in Figure 3. As shown in Figure 6, Lmk-based methods refers to Aniportrait. It is not capable of generating head movement like spinning and nodding since the projected 2D landmarks have lost most of the 3D pose information, thus offering vague pose information in denoising process. In Figure 6, the head generated by Aniportait in the first 3 columns from left are all squished due to this projection error. In the remaining columns, aniportrait failed to generate accurate facial dynamics when there are significant differences between the driving identity and reference identity. As we disscussed in former sections, these are the natural shortcomings of landmarks serving as control signals. However, in audio-driven mode, Aniportrait hardly generate head movement or dramatic facial expressions from audio, making it easier to maintain visual quality than other methods. On the other hand, our model use unified FLAME coefficients as conditions, avoiding the need for alignment and thus generating accurate facial dynamics and head pose as shown in Figure 6. Our approach demonstrates that FLAME coefficients can capture more subtle expressions, and is a more suitable as intermediate representation for controllable generation of talking head avatars.
@@ 6 Conclusion
Figure 1: Additional Results on Multi-pose Generation